OpenStudy (anonymous):

solve (x-3)/(x-4)>equalto 2

OpenStudy (freckles):

\[\frac{x-3}{x-4}-2 \ge 0\] first step I would do is combine fractions on the left

OpenStudy (princeharryyy):

do u need answer or the solution @IloveHW

OpenStudy (freckles):

then find when that fraction is zero and undefined then draw a number line then test the intervals around each number that gave us the fraction was undefined or zero

OpenStudy (amistre64):

-3 = -4+1 just a stray thought tho

OpenStudy (princeharryyy):

x belongs to (4,5]

OpenStudy (amistre64):

1 + 1/(x-4) >= 2 1/(x-4) >= 1

OpenStudy (amistre64):

in order for this to be true, the bottom has to be positive soo the sign can stay the same direction and 1 >= x-4

OpenStudy (anonymous):

can someone take me step by step i dont just want the answer

OpenStudy (amistre64):

tada!!

OpenStudy (princeharryyy):

ohk @IloveHW

OpenStudy (campbell_st):

multiply both sides by the square of the denominator \[\frac{(x - 3)(x - 4)^2}{(x - 4)} \ge 2(x - 4)^2\] remembering \[x \neq 4\] then solve for x

OpenStudy (anonymous):

its not square btw

OpenStudy (princeharryyy):

this is wrong @campbell_st never do it like this

OpenStudy (princeharryyy):

wrong process.. let me help u out @IloveHW

OpenStudy (freckles):

(x-4)^2 is positive I don't see why he can't do it that way

OpenStudy (freckles):

he multiplied both sides by a positive number which doesn't effect the direction of the inequality

OpenStudy (campbell_st):

I know, you take the denominator and square it.... then multiply both sides of the inequality by that

OpenStudy (campbell_st):

it gives you \[x^2 - 7x + 12 \ge 2x^2 - 16x + 32\] next collect like terms... etc

OpenStudy (anonymous):

the solution is x|4<x<eqto5

OpenStudy (freckles):

\[\frac{x-3}{x-4} -2 \ge 0 \\ \frac{x-3}{x-4}-\frac{2(x-4)}{x-4} \ge 0 \\ \frac{x-3-2(x-4)}{x-4} \ge 0\] find when bottom is zero and when top is zero then put those numbers on a number line and test the intervals around those numbers to see which are positive or not

OpenStudy (princeharryyy):